How do you find intercepts of a line in an equation such as y = -3x - 20
Answer by HyperBrain(694)

y=-3x-20
to find the x-intercept/s, set y=0:
-3x-20=0
-3x=20
x=-20/3
to find the y-intercept, set x=0;
y=-3(0)-20=-20
Power up,
HyperBrain!

Help us promote our advocacy:
http://uselessplastic.wordpress.com/
http://uselessplastic.multiply.com/
Post a comment